OpenStudy (rainbow_dash):

Find the sum of the given infinite geometric series 3+1.8+1.08+...

OpenStudy (rainbow_dash):

i knw the formula is \[S \infty = \frac{ a_1 }{ 1-r }\] but don't know what r is or how to find it.

hartnn (hartnn):

r is the common ratio between the terms, like r= 2nd term/1st term = 3rd term / 2nd term so, what will be r here ?

OpenStudy (rainbow_dash):

.6

hartnn (hartnn):

is 1.8 / 3 = 1.08 /1.8 = 0.6 ? if yes, then r= 0.6 is correct

OpenStudy (rainbow_dash):

okay so just 3/.4 then?

OpenStudy (rainbow_dash):

got it, thanks :)

hartnn (hartnn):

yup. thats it

hartnn (hartnn):

welcome ^-^ go through this when u have time .. Tutorial on AP,GP,HP. http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0